Fermi National Laboratory
 

Accelerator Update for February 8, 2006 - February 10, 2006

 

Wednesday February 8

The day shift began with Operations monitoring store 4631 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with experts working on the NuMI beam problems, and with MiniBooNE and the Meson MTest and MCenter experiments taking beam.

Operations began preparing the Recycler to stash antiprotons at 8:40 AM.

At 9:11 AM, MiniBooNE tripped off for the first of many, many times due to a power supply’s (HV865) wandering output. It reset each time. An expert later reported that the trips were not due to the power supply, but to noise.

The Recycler ended its stash. MTest and MCenter resumed taking beam at 9:26 AM. Pbar resumed stacking a few minutes later.

Operations began preparing the Recycler to stash at 12:02 PM.

A controls expert working on the MiniBooNE noise problem reported at 12:27 PM that he had found the cause and was working on a fix.

The stash ended. MTest and MCenter resumed taking beam at 1:37 PM.

At 5:20 PM, a Booster power supply (B:DOGL3) tripped off for the first of several times over the next two hours. An expert investigated twice, but found nothing wrong.

Operations resumed sending beam to NuMI at 6:24 PM.

At 7:31 PM, NuMI tripped off due to a beam problem. Experts investigated and, at 8:07 PM, Operations resumed sending them beam.

Tevatron (TeV) experts reported at 8:22 PM that they were beginning an end-of-store study.

The Recycler began preparing to stash at 8:52 PM.

Operations terminated store 4631 at 9:15 PM.

Operations prepared the D-Zero (D0) collision hall for access at 9:25 PM.

The stash ended at 9:27 PM. Operations turned off beam to Pbar so that operators could inspect the tunnels for a possible LCW leak.

D0 came out of access at 9:59 PM.

Operations put the TeV through a dry squeeze at 10:17 PM.

At 11:13 PM, Operations reported finding a Pbar 95 LCW leak on a coupling for magnet IQ5. An operator called in two LCW technicians.

Operations resumed sending beam to NuMI, MiniBooNE, MCenter, and MTest at 11:37 PM.

Thursday February 9

The midnight shift began with Operations monitoring the stash, and with Pbar off due to an AP1 transfer line LCW leak, which halted beam in the MI-8 transfer line due to its proximity.

The MTest experiment finished using the beam at 12:06 AM. Operations turned their beam off.

The Recycler began preparing to stash at 12:25 AM.

Operations reported that the Water techs had the AP1 LCW leak repaired by 12:46 AM.

MiniBooNE and NuMI resumed taking beam by 12:51 AM.

Pbar resumed stacking at 1:42 AM.

Operations put the TeV into shot setup at 1:52 AM.

Operations established store 4634 at 3:47 AM with an initial luminosity of 162.92E30. Pbar resumed stacking, followed by beam to MCenter and MTest.

At 5:34 AM, Operations reported a network problem; all but two of the MCR consoles were reporting “No Response.” Operations informed experts and then took the master switch, which shut down all beam except for the store and the stack.

A controls expert talked operators through the replacement of a network switch at 6:17 AM. This solved the problem. With the return of the consoles operations resumed normal running.

MCenter began work on their analyzing magnets at 7:13 AM. Operations turned off their beam.

The Recycler began preparing to stash at 10:37 AM.

MCenter resumed taking bean at 2:16 PM.

Pbar resumed stacking around 2:54 PM.

Operations turned off beam to NuMI for work at 3:51 PM. The work was completed and NuMI resumed taking beam at 4:35 PM.

NuMI tripped of at 5 PM. Operations investigated.

Pbar experts began a study period at 6:40 PM. The study ended at 7:41 PM.

An MCenter expert reported at 8:45 PM that they have a LCW leak on their analyzing magnet. They want to run without the magnet, but they have a low LCW level. Operations called in a LCW expert.

NuMI resumed taking beam at 9:13 PM.

The recycler began preparing to stash at 9:19 PM. Operation reported a problem with the first transfer of antiprotons. Operators stopped the stash and investigated.

Operations prepared the Recycler to resume stashing at 11:23 PM.

Friday February 10

The midnight shift began with Operations monitoring store 4634, with the Recycler stashing, and with NuMI and MiniBooNE taking beam.

The stash ended and Pbar resumed stacking at 12:21 AM.

The Recycler prepared to stash at 3:49 AM. The stash ended and Pbar resumed stacking at 4:31 AM.

At 4:57 AM, Operations reported that the Main Injector (MI) Anode power supply tripped off. It would reset, but then trip again. Operators investigated and found a problem with MIRF7. They bypassed the station and MI resumed normal running.

A LCW expert reported at 7:43 AM that the MS6 LCW system was up and running normally.

The Plans for Friday and the Weekend

The plans for today and the weekend are to stack, stash, and store. The Run Coordinator will allow a D0 access and a TeV equipment test in between stores this afternoon.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 02/10/2006    email Fermilab